If 38.2 ml of 0.248M aluminum sulfate solution is diluted with deionized water to a total volume of 0.639L; how many grams of aluminum ion are present in the diluted solution?
Answer #1

intial                                                                    final

M1 = 0.248M                                                     M2 =

V1 = 38.2ml                                                      V2 = 0.639L   = 639ml

            M1V1   =    M2V2

            M2      = M1V1/V2  

                       = 0.248*38.2/639   = 0.0148M

no of moles of Al2(SO4)3   = molarity * volume in L

                                            = 0.0148*0.639   = 0.009457moles

mass of Al2(SO4)3   = no of moles * gram molar mass

                                 = 0.009457*342   = 3.23g >>>>answer